The sum of two consecutive integers plus 7 is 70. Find the integers.
My work so far:
70-7=63
63/2=31.5
c+i=31.5

Answer by josgarithmetic(39630)

Any two consecutive integers may be n and n+1.
(n)+(n+1)+7=70
n+n+1+7=70
2n+8=70
n+4=35
n=35-4
n=31 and n+1=32